OpenStudy (anonymous):

2(3x-3)+5=2(3x-1) show work please?

jagr2713 (jagr2713):

Hey What do you think we do first?

OpenStudy (mokana219):

so first of all you have to use the distributive property, multiply anything right outside the parenthesis into everything inside the parenthesis, so if we distributed both 2s we'd wind up with 6x-6+5=6x-2, you simplify that to 6x-1=6x-2 then solve for x

OpenStudy (moazzam07):

2(3x-3)+5=2(3x-1) 6x-6+5=6x-2 now you get all X on 1 side so then you would subtract 6 from both sides. 6x - 6x = 0 so in that case X=0